Shift Summary for Owl shift, 09/30/99

ugur@cuhep.Colorado.EDU
Thu, 30 Sep 1999 05:56:23 -0600 (MDT)

Shift Summary for Owl shift, 09/29/99

On Shift: L. Bellanto and U. Erkarslan

Hours of beam delivered at all intensities:8.2
Hours of beam delivered at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...):

Accounting of lost time: Quench at A11, Front end problems, Bad beam


========================================================================
Runs ended this shift:

Run 14768 - PEDRUN
Run 14769 - BEAM_799 2.6M
Run 14770 - BEAM_799 Nothing useful
Run 14771 - BEAM_799 Short
Run 14772 - BEAM_799

Tapes brought to FCC this shift:
(none)
========================================================================
Accesses during shift:

none.
========================================================================
Overview of shift:

Pdestal run. Difficult E799 datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:00 Take over the shift, no beam, cosmic run

00:15 Wah and Leo notice a massive dip in DC3x hits in
cosmic configuration - Talked w/ Hogan, decided to
leave it. It disappeared with beam
Decide to skip the muon run -haven't gotten that mud
data since last one
14768 00:19 Pedestal run started-Magnet still ramping, Beam is
back.
FPED, CsI ped update requested. We just gave FPED
update.
01:10 Got many scaler error. Phoned MCR. They say they are
tunning extraction. DC's are ok but drawing current-
Ask for beamstop in till they are done.

Spill signal absent at scope for monitoring.
Investgation shows is not clear that the
429 on left end of 2nd NIM rack in RR160 has any
functional outputs left-- but data taking seems OK
otherwise
01:25 Trgchk is not update since 12:52. Paged Theo. Can not
find trgchk on sub_job.tcl. Ultimate reason was that
this is a SERVER, not a PROCESS and one must try to
select the start SERVER options in sub_job
01:38 FPED stuck; KFREND fixed it
01:49 Spill sounds smoother. Phone MCR. They think the
extraction QXR will learn itself in the rest of the
way. Ask for them to open beam stops
14769 55 01:51 Beam stops open. Spill still crappy. DC current not so
bad. Spill uniformity bad throughout shift
66 Unusual error from Pipline asks for reinit after print
out of file on log book XXI page32
$pipe_DIR/drc/status/ktstat_03.
14769 85 02:35 PMT gain drift wrt dbase looks bad; long tail on
negative side. Suspecting that this histogram could
have been affected by the ~60 spills taken with
beamstops in, tried to reset histograms - does not
seem to reset this histogram.
14769 97 02:48 Try stopping and starting a new run but fail.
Stream 5 top DCY goes nonempty. Try drt_fresh_start.
ultimately solved by cycling downstairs repeater
a few moments before ...
03:31 Quench at A11
14771 71 04:50 Beam back - no laser monitoring histo at all
05:08 Try stopping and restart with brute force on all
options. Still no CsI monitoring. Decide to run
blind for the last 2 hours.
06:25 SDAQ red button.Scaler Check is not updated. Same
problem we think with not being able to restart
servers.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Examine 429 on far left of 2nd NIM crate in RR160 (trigger racks) to see
if it is really working; if so, get spill signal to scope in
control room.

Find out why we seem to have lost the laser monitoring

Take a laser scan

Track down unusual pipeline error

Flip NM2SR. Nobody has done this for weeks I think.

E/p is a little wide, at sigma 1.56; RF timing at 94.5ns creates frownies

SOD width on 3X is a little wide. Wonder if it is related to the dead
channels seen in the cosmics running.~

All the things on Bob T's list for the shutdown; I think that tracking
down the source of the RC10 scaler errors is on this list

Keep an eye peeled for the new Xerox machine. Should arrive in time for
KaMI